# Automated Defects Detection and Fix in Logging Statement

### v1

@@ 6. Threats to Validity
Human involvement. We conduct manual studies to identify defects in logging statements and human evaluation of new repositories. This human-involved process may introduce bias during annotation. To minimize this threat, we engage two independent annotators with expertise in the field, and ensure that any discrepancies are resolved through discussion until a consensus is reached. Besides, we get 0.78 Cohen’s Kappa for manual defeat identification and 0.84 for human evaluation, which is a substantial level of agreement.
@@ 7. Related Work
Studies on maintaining high-quality logging statement. Many related works focused on potential issues associated with logging statements. Li et al. (Li et al., 2023a) focused on the readability of log messages in the logging statements, they derived three aspects related to the readability and explored the possibility of automatically classifying the readability of log messages. Ding et al. (Ding et al., 2023) made the first attempt to comprehensively study the temporal relations between logging statements and corresponding source codes. They proposed a tool to detect the issues of temporal inconsistencies. Bouzenia et al. (Bouzenia and Pradel, 2023) targeted checking for inconsistencies between conditions and logging statements and presented CMI-Finder, a neural network-based model to detect condition-statement inconsistencies. Chen et al. (Chen and Jiang, 2019, 2017) characterized and detected the anti-patterns in the logging statements. They encoded six anti-patterns into a static code analysis tool to detect previously unknown anti-patterns in the source code.
@@
The primary distinction between our work and the aforementioned studies lies in the fact that while these studies solely detected specific categories that may affect the quality of logging statements, our approach not only identifies these issues but also provides solutions for their rectification.
@@
Empirical studies on the existing logging practices. Owing to advancements in logging within software engineering, there has been a surge in research exploring logging practices. Li et al. (Li et al., 2021b) conducted an in-depth qualitative analysis of both the benefits and limitations of logging in software development. Zhou et al. (Zhou et al., 2020) delved into the ramifications of logging practices on data leaks within mobile apps. Kabinna et al. (Kabinna et al., 2018) identified that aspects like bug corrections, feature augmentations, and code restructuring often led to modifications in logging statements. Recently, Li et al. (Li et al., 2023b) pioneered a study on LLMs for logging statement generation and discovered that employing prompt-based strategies with zero-shot or few-shot learning enhances the generalization capabilities of LLMs.
8. Conclusion
In this paper, we first conduct a pilot study and identify four types of logging statement defects that developers are concerned about, which can affect the quality of runtime logs. Then we propose LogFixer, a framework for automatically detecting and updating potential defective logging statements. Experiment results demonstrate the effective detection and updating capabilities and practicality on new projects.

### v2

@@ 6.2.2. Internal validity
Potential data leakage. We conduct experiments using various LLM backbones. Because the training data for these models are undisclosed, there is a risk of potential data leakage. To mitigate this, we initiate our data collection in April 2024, postdating the training of most LLMs. However, we recognize that for the existing logpatches dataset, there may be some risk of data leakage when testing the updating effects. To mitigate data leakage, we add synthetic test data and new repositories data. The synthetic test data are generated by the proposed synthetic strategies, which can ensure that LLMs cannot include these data as part of their training dataset. For new repositories, code snippets are current as of April 2024. The LLM updates defective logging statements without relying on April 2024 information, reducing data leakage risk. Experimental results demonstrate that LogUpdater can achieve a 61.49% success rate with new repository data, proving our method’s effectiveness. Human involvement. We conduct manual studies to identify defects in logging statements and human evaluation of new repositories. This human-involved process may introduce bias during annotation. To minimize this threat, we engage two independent annotators with expertise in the field, and ensure that any discrepancies are resolved through discussion until a consensus is reached. Besides, we get 0.78 Cohen’s Kappa for manual defect identification and 0.84 for human evaluation, which is a substantial level of agreement.
6.2.3. External validity
The selection of programming languages. We source data from 641 Java projects, which may impact the generalizability of our method for other languages. Following previous study (Ding et al., 2023), we choose Java because it is a mature language with multiple built-in logging and third-party logging frameworks. Nevertheless, applying LogUpdater to other programming languages may affect its performance because of distinct language features. While fine-tuning detectors on other language-specific datasets is a promising solution, we aim to incorporate more languages in future work. The selection of LLM backbone. We select GPT4o as our basic LLM backbone and employ six different popular instruction-taken LLMs for experiments to demonstrate the effectiveness of LogUpdater. However, different LLM models and model versions exhibit varying levels of comprehension regarding prompts, leading to fluctuations in model performance during experiments. To ensure reproducibility, we set the temperature to 0, theoretically allowing the model to return the same results for identical queries. Additionally, we plan to extend the relevant experiments to newer emerging LLMs to further assess the generalizability of our approach.
@@ 7. Related Work
7.1. Studies on Maintaining High-quality Logging Statement
In numerous related studies, researchers have concentrated on identifying potential problems associated with logging statements. Li et al. (Li et al., 2023a) dedicated their research to analyzing the readability of log messages within these logging statements, formulating three dimensions pertinent to readability. They delved into the feasibility of implementing automatic classification systems for the readability of log messages. Conversely, Ding et al. (Ding et al., 2023) embarked on a pioneering investigation into the temporal relationship between logging statements and their corresponding source code. Their efforts culminated in the development of a tool designed to pinpoint and address temporal inconsistencies. In a different vein, Bouzenia et al. (Bouzenia and Pradel, 2023) directed their focus towards detecting discrepancies between conditions and logging statements, unveiling CMI-Finder—a model grounded in neural networks capable of identifying inconsistencies in condition-statement pairs. Zhang et al. (Zhang et al., 2022) undertook a comprehensive analysis of 21 open-source projects, encompassing 70,000 logging statements, which include 48,000 production logging statements and 22,000 test logging statements. This study provided developers and researchers with substantive insights into the interplay between test and production logging. Chen et al. (Chen and Jiang, 2019, 2017a) explored and pinpointed anti-patterns manifested in logging statements, incorporating six of these anti-patterns into a static code analysis tool to unearth previously undetected anti-patterns in source code. Additionally, Li et al. (Li et al., 2019, 2021c) pursued an in-depth investigation into duplicate logging statements. By meticulously examining over 3,000 duplicate logging statements and their corresponding code across four expansive systems, they delineated five patterns indicative of duplicate logging smells. Furthermore, they introduced DLFinder, a tool engineered to automatically identify problematic duplicate logging code smells.
@@ 7.1. Studies on Maintaining High-quality Logging Statement
The primary distinction between our work and the aforementioned studies lies in the fact that while these studies solely detected specific categories, our approach not only identifies these issues but also provides solutions for their rectification.
7.2. Empirical Studies on Existing Logging Practices
@@
Owing to advancements in logging within software engineering, there has been a surge in research exploring logging practices. Li et al. (Li et al., 2021b) conducted an in-depth qualitative analysis of both the benefits and limitations of logging in software development. Zhou et al. (Zhou et al., 2020) delved into the ramifications of logging practices on data leaks within mobile apps. Kabinna et al. (Kabinna et al., 2018) identified that aspects like bug corrections, feature augmentations, and code restructuring often led to modifications in logging statements. Recently, Li et al. (Li et al., 2023b) pioneered a study on LLMs for logging statement generation and discovered that employing prompt-based strategies with zero-shot or few-shot learning enhances the generalization capabilities of LLMs. Zhao et al. (Zhao et al., 2023) conducted an empirical study on the IDs in logging statements, proposed a simple approach to inject IDs in logging statements to mitigate information loss issues, and studied the information gained by injecting IDs into logs. Zeng et al. (Zeng et al., 2019) and Chen (Chen and Jiang, 2017a) extended the study of Yuan et al. (Yuan et al., 2012a) to Android and Java systems, finding a massive presence of log statements in the analyzed systems. Li et al. (Li et al., 2023c) investigate the characteristics of dynamic variables and their importance in practice, and then explore the potential of a variable-aware log abstraction technique. Lai et al. (Lal et al., 2015) perform an analysis of logging code constructs at two levels. They answer nine research questions related to statistical and content analysis at file level and block level. Additionally, significant studies have been made in determining log levels (Liu et al., 2022; Li et al., 2021a; Tang et al., 2022) and content (Ding et al., 2022; Liu et al., 2019).
7.3. Logging Statement Automation
Conventionally, the process of logging statement automation can be categorized into two stages (Chen and Jiang, 2021; He et al., 2021): identifying logging locations and creating logging statements. We refer to these steps as where-to-log and what-to-log (Li et al., 2024a). To address the challenge of determining where to log, researchers have explored various methods (Jia et al., 2018; Li et al., 2020; Yao et al., 2018; Zhao et al., 2017; Zhu et al., 2015; Yuan et al., 2012b; Li et al., 2018). to identify suitable logging locations within the source code. In terms of what to log, the generation of logging statements is typically broken down into three subtasks: generating logging text (Ding et al., 2022; Mastropaolo et al., 2022; Xu et al., 2024; Xie et al., 2024), selecting logging variables (Liu et al., 2019; Yuan et al., 2012d; Dai et al., 2022), and predicting the logging level (Heng et al., 2024; Li et al., 2017, 2021a; Liu et al., 2022; Mizouchi et al., 2019). The most recent approach (Unilog (Xu et al., 2024), Fastlog (Xie et al., 2024), SCLogger (Li et al., 2024a), LANCE (Mastropaolo et al., 2022)) provides automatically logging statements generation solution of deciding logging places, statements level, content, and variables at one-step by leveraging the ability of LLMs.
As we detailed discussion in Section 6.1.2, although recent works have investigated the capabilities of LLMs in generating logging statements automatically, LogUpdater continues to offer unique benefits by lower costs and better performance in detecting and updating defective logging statements.
8. Conclusion and Future Work
In this paper, we find that existing research on analyzing existing logging statements is limited, primarily focusing on detecting a singular type of defect (limited scope of analysis) and relying on manual intervention for fixes rather than automated solutions (detection without repairing).
First, we conduct a broad-scope pilot study and identify four types of common defects that developers are concerned about from 641 targeted repositories. This coverage enables us to analyze defective logging practices across varied development environments, unveiling new insights into developer priorities and defect patterns that are more broadly representative of real-world software development.
Then, we propose a two-phase end-to-end framework to automatically detect and update potential defective logging statements. This is a pioneering advancement in the field, especially valuable for large-scale projects where manual repair is not feasible. By automating repairs, our tool saves significant manual effort, addressing a gap not covered by existing literature. Extensive experiment results demonstrate the effective detection and updating capabilities and practicality of LogUpdater.
In future work, we will extend LogUpdater to handle log-level optimization using context-aware techniques, addressing project-specific conventions while maintaining the framework’s reliability. Specifically, a potential path is to develop a context-aware detection framework combining organizational guidelines and runtime execution frequency. By learning cross-project commit histories and logging statements, the model will learn optimal level assignments that balance observability needs with overhead constraints. The framework will adaptively detect misuse log-level based on code context, enabling project-specific yet automated existing logging statement checking.
Acknowledgement
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).